Merger Analysis Textbook answer is: Vop unleavered = $32.02 million; Vtax sheilds = $11.50...

Marston Marble Corporation is considering a merger with the Conroy Concrete Company. Conroy is a publicly traded company, and its beta is 1.30. Conroy has been barelyprofitable, so it has paid an average of only 20% in taxes during the last
several years. In addition, it uses little debt; its target ratio is just 25%, with the cost of debt 9%.
If the acquisition were made, Marston would operate Conroy as a separate, wholly owned subsidiary. Marston would pay taxes on a consolidated basis, and the tax ratewould therefore increase to 35%. Marston also would increase the debt capitalization in the Conroy subsidiary to wd = 40%, for a total of $22.27 million in debt by theend of Year 4, and pay 9.5% on the debt. Marston’s acquisition department estimates that Conroy, if acquired, would generate the following free cash flows and interestexpenses (in millions of dollars) in Years 1–5:





Year Free Cash Flows Interest Expense
1 $1.30 $1.2
2 1.50 1.7
3 1.75 2.8
4 2.00 2.1
5 2.12 ?


In Year 5, Conroy’s interest expense would be based on its beginning-of-year (that is, the end-of-Year-4) debt, and in subsequent years both interest expense and freecash flows are projected to grow at a rate of 6%.
These cash flows include all acquisition effects. Marston’s cost of equity is 10.5%, its beta is 1.0, and its cost of debt is 9.5%. The risk-free rate is 6%, and themarket risk premium is 4.5%.
a. What is the value of Conroy’s unlevered operations, and what is the value of Conroy’s tax shields under the proposed merger and financing arrangements?
Answer #1

a) rSL = 5 + 1.0*3.5 = 8.5% rSU = 0.30*9+0.70*8.5 = 8.65% 8.5 = rSU + (rSU - rD)*30/70; 8.5 = rSU + (rSU-9)*3/7; 8.5 = (10/7)*rSU - 27/7; 12.3571 = 10/7*rSU; rSU = 8.65 % Value of unlevered operations: 1 2 3 4 5 FCF ($ million) 1300000 1500000 1750000 2000000 2120000 pvif at 8.65% 0.92039 0.84711 0.77967 0.71760 0.66047 PV of Horizon FCF 1196503 1270667 1364422 1435196 1400191 CUM PV of Horizon FCF 6666978 Terminal value of FCF = 2120000*1.06/(0.0865-0.06) 84800000 PV of terminal FCF=84800000*0.66047 = 56007630 Intrinsic unlevered value of operations 62674609 = $62.67 millions


answered by: Bijay Agrawal